### v1

# CoCoEvo: Co-Evolution of Programs and Test Cases to Enhance Code Generation

Abstract
Large Language Models (LLMs) have shown remarkable performance in automated code generation. However, existing approaches often rely heavily on pre-defined test cases, which become impractical in scenarios where such cases are unavailable. While prior works explore filtering techniques between programs and test cases, they overlook the refinement of test cases. To address this limitation, we introduce CoCoEvo, a novel LLM-based co-evolution framework that simultaneously evolves programs and test cases. CoCoEvo eliminates the dependency on pre-defined test cases by generating both programs and test cases directly from natural language problem descriptions and function headers. The framework employs specialized evolutionary operators, including LLM-based crossover and mutation operators for program evolution, along with a test case generation operator for test case evolution. Additionally, we propose optimization strategies such as a crossover rate scheduler to balance exploration and convergence, and a multi-objective optimization method for test case selection. Experimental results on multiple state-of-the-art LLMs demonstrate that CoCoEvo surpasses existing methods, achieving state-of-the-art performance in automated code generation and testing. These results underscore the potential of co-evolutionary techniques in advancing the field of automated programming.
Index Terms:
Large Language Models, Code Generation, Test Case Generation, Co-EvolutionI Introduction
In recent years, Large Language Models (LLMs) have undergone rapid development and have been widely applied in automated software development tasks such as code generation and test case generation. Models like GPT-4 [1], Llama3 [2], Qwen2.5 [3], and DeepSeek-V3 [4] have demonstrated exceptional code generation capabilities, significantly advancing the field of automated software development.
Despite their impressive performance, LLMs do not always generate accurate or correct programs. To enhance the reliability of program generation, test cases are often utilized to evaluate code correctness and provide feedback. Existing approaches frequently rely on pre-defined test cases for evaluation. For instance, Sampling+Filtering [5] first generates a batch of candidate codes, evaluates them using pre-defined test cases, and selects the code with the highest score as the final solution. Other methods, such as Reflexion [6], Self-Repair [7], INTERVENOR [8], LDB [9], and PairCoder [10], utilize execution feedback on pre-defined test cases to refine or repair generated code. These methods generally assume that pre-defined test cases are accurate and can be fully trusted.
However, pre-defined test cases are not always available. In test-driven development (TDD) [11], test cases are defined before program implementation. In real-world scenarios, software development frequently relies on natural language requirements, with no pre-existing test cases or example programs. Furthermore, creating pre-defined test cases can be labor-intensive, requiring significant manual effort. In situations where LLMs must generate test cases independently, the effectiveness of existing methods remains uncertain.
Some approaches do not rely on pre-defined test cases. For instance, methods like CodeT [12] and MBR-Exec [13] generate both code and test cases using LLMs and filter the code based on the execution results of these test cases. Similarly, CodeCoT [14], which utilizes chain-of-thought prompting [15], enables LLMs to generate code and test cases simultaneously and evaluate and refine the code based on these cases. Additionally, AgentCoder [16] introduces a multi-agent system with a dedicated Tester agent to generate test cases for evaluating and repairing code. However, studies such as [12, 17] highlight that LLM-generated test cases can contain significant errors, potentially leading to unreliable evaluations and incorrect feedback. Moreover, many of these methods neglect to verify the quality of the generated test cases. While approaches like CodeT and MBR-Exec implement basic filtering mechanisms for test cases, other methods often use these generated test cases directly for code evaluation without ensuring their correctness.
There are also previous co-evolution methods based on code and test cases, which have primarily focused on code maintenance and program repair. For example, [18] recommends new test cases based on source code similarity, [19] identifies outdated test cases, and [20] constructs test oracles from bug reports using predefined templates. Similarly, [21] generates test cases based on predefined specifications to drive software evolution. However, these approaches depend heavily on pre-existing programs, test specifications, or templates, limiting their applicability in scenarios lacking such resources.
To address these limitations, we propose CoCoEvo, an LLM-based co-evolution framework that enables the simultaneous evolution of programs and test cases to produce more accurate outcomes (illustrated in Fig. 1). In CoCoEvo, both test cases and programs are generated by LLMs, requiring only natural language problem descriptions and function headers. No pre-defined programs, test cases, or templates are needed. The CoCoEvo framework consists of two alternating steps: program evolution and test case evolution. In program evolution, we introduce LLM-based crossover and mutation operators to generate program offspring. The current population of test cases is used to evaluate program fitness. In test case evolution, we design an additional test case generation operator to produce test case offspring. Conversely, the current population of programs is used to evaluate test case fitness. For programs, we implement a fitness calculation method inspired by CodeT [12], which assesses agreement between programs and test cases. For test cases, we adopt a multi-objective optimization approach to select those with high accuracy and strong discriminatory power, using the Pareto method [22]. Additionally, we introduce a dynamic crossover rate scheduling mechanism, enabling broader exploration during early evolution stages and faster convergence later.
To evaluate the effectiveness of CoCoEvo, we collected a LeetCode-Contest dataset comprising relatively new programming problems that are unlikely to have been included in the training data of the evaluated LLMs. We conducted experiments on four state-of-the-art models: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4]. We compared with both basic code generation methods such as Sampling [5] and Sampling+Filtering [5], repair-based methods, and agreement-based methods. The results demonstrate that CoCoEvo outperforms all existing methods. Additionally, the experiments revealed that methods relying on pre-defined test cases for feedback experience significant performance degradation when adapted to self-generated test cases. This highlights the limitations of such methods in scenarios lacking pre-defined test cases. Furthermore, we conducted in-depth comparative experiments and discussions, which validated the effectiveness of the optimization strategies employed in CoCoEvo.
In summary, our main contributions are as follows:
1.
Co-evolution framework. We propose a novel LLM-based co-evolution framework for programs and test cases, including specialized LLM-based operators for offspring generation.
2.
Optimized Evolution Techniques. For program evolution, we design a dynamic crossover rate scheduler that enhances exploration early on and accelerates convergence later. For test cases, we develop a multi-objective optimization method for fitness evaluation and selection.
3.
Experimental Validation. We introduce the LeetCode-Contest dataset and conduct experiments using four state-of-the-art LLMs: GPT-4o-mini, Qwen2.5-Coder-32b, Llama-3.1-70B, and DeepSeek-V3. The experimental results demonstrate that CoCoEvo achieves state-of-the-art performance, surpassing existing methods. Additionally, we conduct ablation studies to evaluate the utility of each module within the framework, further highlighting the effectiveness of our approach.
The remainder of this article is organized as follows. Section II discusses the related work. In Section III, the design details and implementation methods of the CoCoEvo are described. Sections IV and V present the experimental settings and discussions, respectively. Finally, Section VI concludes this article and outlines future work.
II Related Work
II-A Enhancing Code Generation Accuracy with Test Case
The integration of test cases has emerged as a pivotal strategy for improving the quality of code generated by LLMs. Numerous studies leverage test cases to filter, evaluate, and refine generated programs, leading to higher-performing and more reliable code. Below, we provide an overview of notable methodologies in this domain.
Li et al. [5] introduce a Sampling+Filtering approach, where a large set of candidate programs is sampled and scored based on performance against a provided test set. Chen et al. [12] propose a framework in which both code and test cases are generated by LLMs simultaneously, using a dual agreement metric to identify and select better programs. Shi et al. [13] assess semantic similarity by executing generated programs on test cases and leveraging the execution results to guide program selection.
Beyond evaluation, test cases play a critical role in iterative repair and refinement of generated code. Execution outcomes from test cases can be used to repair faulty programs. For instance, Shinn et al. [6] introduce the Reflexion framework, where failure messages from test executions are converted into verbal feedback for program repair. Similarly, Olausson et al. [7] evaluate LLMs’ capabilities to repair programs using failed test feedback. Zhang et al. [23] allow LLMs to analyze test execution results to revise erroneous code. Zhong et al. [9] extend this idea by enabling step-by-step debugging. Huang et al. [14] adopt a chain-of-thought prompting [15] approach that generates code and tests sequentially, iteratively improving the code based on test feedback. Zhang et al. [10] introduce a pair programming framework where test feedback informs repair strategies and planning.
The generation of test cases differs across these methodologies. For example, CodeT [12] and CodeCoT [14] rely entirely on LLMs to generate test cases, whereas Reflexion employs LLMs primarily for simple programming tasks. Most other methods, however, assume pre-defined test cases as part of their evaluation frameworks, leveraging them for filtering, refinement, and program selection.
II-B Automated Test Case Generation
Early works, such as [24], leveraged search-based methods to generate test cases. More recently, there has been a growing trend toward utilizing LLMs for test case generation. Studies such as [25], [26], and [27] have explored the effectiveness of LLMs in automating program testing. Similarly, [28] investigates methods for enhancing existing test cases using LLMs. Other approaches, such as TOGA [29], TOGLL [30], and the method described in [31], focus on generating test oracles based on predefined templates. Meanwhile, [32] introduces a model trained with deep reinforcement learning for text-to-test case generation. The application of LLMs in software development extends to generating test cases for various scenarios. For instance, [33] employs GPT-4 to generate security tests for identifying vulnerabilities, while [27] uses LLMs to assist in generating unit tests for automatic software testing. Similarly, [34] demonstrates the use of LLMs to create test cases aimed at reproducing general software bugs. LLMs are also integrated into code generation studies to improve the accuracy of generated code through test case generation. CodeT [12], for example, utilizes LLMs with zero-shot prompts to directly generate test cases. Reflexion [6] introduces a Test Agent powered by LLMs for test case generation.
II-C Evolution with LLMs for Code Generation
The integration of LLMs with evolutionary algorithms has been extensively explored to enhance code generation. For example, Bradley et al. [35] introduced OpenELM, an open-source Python library that designs specialized evolutionary operators for code generation. Chen et al. [36] utilized LLMs as adaptive mutation and crossover operators in evolutionary neural architecture search (NAS), optimizing neural network design by merging LLM capabilities with evolutionary processes. Liu et al. [37] proposed an automated approach for evolving optimization algorithms using LLMs, which eliminates the need for manual design by applying standard evolutionary operations such as initialization, selection, and mutation. In a related study, Liu et al. [38] developed a framework for evolving heuristics by integrating LLMs with evolutionary computation to co-evolve both natural language descriptions and executable code for heuristic design, further extending previous methodologies.
Ma et al. [39] introduced an algorithm combining LLMs with evolutionary search to generate and optimize reward functions for reinforcement learning tasks, incorporating a reward reflection mechanism for refinement. Meyerson et al. [40] demonstrated the use of few-shot prompting with LLMs to generate offspring in evolutionary algorithms, enabling a general crossover operator by concatenating parent solutions into a single prompt, which can adapt to various domains. Similarly, Romera et al. [41] integrated LLMs with evolutionary algorithms to create context-aware program mutations, eliminating the need for manually defined mutation operators. Ye et al. [42] proposed ReEvo, which combines evolutionary search with LLM-driven reflections to generate optimization heuristics, efficiently exploring heuristic spaces while offering verbal gradients for optimization. Lastly, AutoTest [43] combines automated test case generation with code execution in an evolutionary framework, utilizing LLMs to co-generate code solutions and their corresponding test cases for enhanced optimization.
Additionally, various LLM-based co-evolution methods have been proposed. For instance, Chi et al. [19] identifies outdated test cases by co-evolving production and test code. Ruan et al. [20] constructs test oracles from bug reports using predefined templates. Chi et al. [19] leveraged LLMs with dynamic validation to automate co-evolution between production and test code, identifying obsolete test cases to improve software quality and reduce maintenance overhead. TestART [44] enhances unit testing for LLMs by combining automated test generation with repair, using template-based fixes and coverage feedback to improve test accuracy.
III Methods
III-A Framework Overview
An overview of CoCoEvo is presented in Fig. 2. The evolutionary process begins with the random initialization of both the program population and the test case population, achieved through the random generating programs and test cases using LLMs. The co-evolution process follows a loop consisting of two alternating steps: program evolution and test case evolution, each of which includes a Generation-Evaluation-Selection sequence. During program evolution, the test case population is used to evaluate programs and calculate their fitness. Conversely, during test case evolution, the program population is utilized to evaluate test cases.
For programs, we introduce LLM-based crossover and LLM-based mutation operators for offspring generation. For test cases, we implement an LLM-based additional test case generation operator. Once the preset number of iterations is reached, the co-evolution loop terminates, and the program with the highest fitness score is returned as the final solution. Algorithm 1 provides the pseudo-code detailing the CoCoEvo process. Further implementation details for each algorithm step are elaborated in subsequent sections.
III-B Program Evolution
This section provides a detailed description of strategies employed in program evolution.
Program fitness function. During evaluation, programs and test cases undergo cross-evaluation, as illustrated in Fig. 3. The outcomes are represented by a matrix :
We compute the confidence score of each program, adopting the approach in [12]. Programs that pass the same set of test cases are grouped into a set , while the corresponding test cases are grouped into a set . The confidence of program is then defined as
where represents the set containing program , and represents the associated test set. The confidence score evaluates the likelihood of a program being correct. The fitness of program is defined as
where a higher fitness score indicates a higher probability of correctness for the program.
Crossover Rate Scheduler To address the multimodal nature of the program search space and mitigate the risk of premature convergence, a dynamic scheduler is employed to regulate crossover and mutation rates. A higher mutation rate is applied in the early stages to promote exploration, while the mutation rate decreases, and the crossover rate increases in later stages to refine convergence.
Unlike traditional random mutation methods, the LLM-based program mutation operator leverages LLMs for code generation, which could retain key features of high-quality programs. This allows for a higher mutation rate without significant loss of desirable traits. We adopt the cosine annealing method [45] for scheduling crossover rate:
This formulation ensures that the crossover rate increases smoothly from to . The number of crossover operations and the number of mutation operations are defined as follows:
where is the program population size, and represent the number of program crossover and mutation operations in iteration round , respectively. When the initialization is completed, the iteration round is set to . After initialization, the iteration round starts at and increments by after a co-evolution loop.
Fig. 4 illustrates the variation in crossover rates over evolutionary rounds. This approach ensures a smooth transition between exploration (early stages) and convergence (later stages), enabling effective optimization of the solution space.
III-C Test Case Evolution
Test fitness function. The evaluation of test cases involves two key metrics: confidence and discrimination. The confidence of test case is calculated as
where represents the total number of programs, represents the fitness of program , and is the evaluation matrix indicating whether passed . Intuitively, confidence measures the degree of agreement between the program population and the test cases, with higher program fitness contributing more weight to the degree of agreement.
Simple test cases may fail to differentiate between correct and erroneous programs, as both types may pass them. To address this, we introduce a discrimination metric. First, the pass rate of test case across the program population is defined as
then the discrimination of test case is represented by the entropy of the pass rate:
Test cases achieve higher discrimination when their pass rate approaches , as they better differentiate among the program population. This indicator assesses whether the test cases can exert evolutionary pressure on the program population.
The selection of test cases is performed using the Pareto approach for multi-objective optimization [22], as illustrated in Fig. 5. The Pareto front is constructed based on the two optimization metrics, confidence and discrimination. Test cases with high confidence and low discrimination represent basic test cases that cover common scenarios, with a high probability of being correct. Test cases with high discrimination represent more challenging cases that are effective at exposing errors. We noticed that there are some test cases with very low confidence levels that sometimes have a very high degree of discrimination. To avoid this type of test case rom being selected, the test cases with confidence lower than the average in the selection results will be filtered out.
III-D LLM-based evolution operators
Program crossover operator.
We utilize LLMs to perform program crossover, as illustrated in Fig. 6. The process involves the following steps:
1.
Selection. Two programs are selected as crossover parents using the binary tournament selection method based on their fitness scores.
2.
Crossover. The selected programs are combined into a prompt. The LLM is instructed to analyze their similarities and differences and generate a new program that merges useful elements from both.
For example, in Fig. 6, two incorrect programs are sent to the LLM. By analyzing their structures, the LLM combines useful fragments from each and generates a new program that resolves the errors.
Program mutation operator. The LLM-based mutation operator intelligently rewrites a program to achieve the same functionality using a different approach. The steps are as follows:
1.
Selection. A program is randomly selected from the current population as the mutation parent.
2.
Mutation. The parent program is included in a prompt, and the LLM is instructed to rewrite it using an alternative implementation method.
For example, Fig. 7 demonstrates a mutation scenario where the task is to find the maximum difference between adjacent elements in a 2D grid. The parent program has a flaw—it only considers elements in the same row or column. The LLM analyzes the problem and rewrites the program to include diagonal relationships, successfully resolving the flaw.
Additional test cases generation operator. LLMs are also employed to generate additional test cases, as shown in Fig. 8. The process includes the following steps:
1.
Information Construct. The prompt contains the following information: The entire existing population of test cases, the best program the program population and the line coverage information on it, where covered lines are marked as “[+]” and uncovered lines as “[-]”.
2.
Rethinking and Generation. Based on the information, the LLMs are required to generate additional test cases: If there are uncovered lines, the LLMs generate test cases to achieve full line coverage. If all lines are already covered, the LLMs analyze the existing test cases to identify untested boundary conditions that could reveal potential program flaws.
For instance, in Fig. 8, the LLM identifies an untested boundary scenario: the input list is not in ascending order. It then generates a new test case with a non-ascending , exposing the flaw in the provided program.
IV Experiments
IV-A Dataset
We collected a dataset of programming problems from the weekly and biweekly contests hosted on the LeetCode platform 111https://leetcode.com/, forming the LeetCode-Contest dataset. In the dataset, each problem includes a prompt comprising the function header and a natural language description in the form of a docstring, along with test cases and reference solutions. To ensure data integrity and avoid potential leakage, all selected problems were released after March 2024.
To more accurately assess the ability of LLMs to solve programming challenges, we took measures to prevent plagiarism of example test inputs and outputs in test case generation. Specifically, we removed all test input-output samples from the problem descriptions, retaining only the natural language problem statements, the function header, and the associated constraints.
For each problem, we collected Python solution codes and a substantial set of test cases available on the website. The detailed information of the LeetCode-Contest dataset is summarized in Table I.
| Problems | 80 |
| Average Test Cases | 644.40 |
| Selected Contests | weekly-contest-402, weekly-contest-401, biweekly-contest-132, weekly-contest-400, weekly-contest-399, biweekly-contest-131, weekly-contest-398, weekly-contest-397, biweekly-contest-130, weekly-contest-396, weekly-contest-395, biweekly-contest-129, weekly-contest-394, weekly-contest-393, biweekly-contest-128, weekly-contest-392, weekly-contest-391, biweekly-contest-127, weekly-contest-390, weekly-contest-389 |
IV-B Models
We utilized four widely adopted and powerful models for our experiments: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4]. For GPT-4o-mini, we accessed the API via the OpenAI platform 222https://platform.openai.com/. For Qwen2.5-Coder-32B and Llama-3.1-70B, we used the APIs available on the DeepInfra platform 333https://deepinfra.com/. For DeepSeek-V3, we utilized the API provided by the DeepSeek platform 444https://platform.deepseek.com/.
IV-C Compared Baselines
We selected several of the most prominent and effective methods in the field of code generation with LLMs as our comparison baselines:
1.
Sampling [5]. A basic method that generates a large number of code samples and randomly selects one as the final answer.
2.
Sampling+Filtering [5]. An enhancement of the Sampling method, where the generated codes are evaluated using test cases. The code with the highest score is selected as the final answer.
3.
Self-Repair [7]. This method generates code randomly, and then feeds execution results on test cases back to the LLM, which repairs the code accordingly.
4.
Reflexion [6]. Error information from test case executions is used to guide the LLM in providing reflection and repair suggestions. The code is then repaired based on these suggestions.
5.
INTERVENOR [8]. Utilizes a Teacher-Student multi-agent framework. The Teacher agent analyzes error information from test case execution and provides repair suggestions, while the Student agent applies the fixes.
- 6.
7.
AgentCoder [16]. Implements two LLM-based agents: a Programming agent for code generation and repair, and a Testing agent for test case generation.
8.
MBR-Exec [13]. Generates both code and test cases using LLMs. A loss function evaluates execution similarity based on the outputs of the generated code when run on test inputs. Codes with higher execution divergence receive higher losses, and the code with the lowest loss is selected as the final answer.
9.
CodeT [12]. Generates code and test cases with LLMs, then performs clustering based on how the codes perform on the test cases. The clustering score is calculated as the product of the number of codes in a cluster and the number of test cases passed.
For all baselines, if the method requires test cases for evaluation or feedback generation, the test cases are generated using the same LLMs that perform code generation.
IV-D Performance Indicators
The ground truth test cases in the dataset are used to evaluate the accuracy of the generated code. A code is considered correct if it passes all the test cases. The performance of the approaches on the dataset is measured using the pass@1 metric [46]. To calculate the pass@1 metric, each method submits a single program as the final answer for each problem. The pass@1 metric represents the percentage of problems successfully solved.
Additionally, for further analysis of the CoCoEvo approach, the ground truth program solutions are used to evaluate the correctness of the generated test cases. A generated test case is considered correct if it executes successfully on the ground truth solution.
IV-E Experimental Settings
Given that invoking LLMs for code generation or repair consumes significantly more time and resources compared to executing code, we evaluated the performance of all methods under the condition that the number of LLM calls for code generation or repair is held constant. To ensure practical usability and align with the parameter settings of most code generation methods, we set the number of LLM calls for code generation to iterations. For methods requiring LLM-generated test cases, we limited test case generation to iterations, with a maximum of test cases extracted per iteration.
The parameter settings of the CoCoEvo method can be found in Table II.
| Program Population Size | 10 |
| Iterations | 10 |
| Min Test Case Population Size | 10 |
We observed that when calculating the pass@1 metric, if multiple codes achieve the same score, one code will be randomly selected as the final result. This introduces a certain degree of randomness to the evaluation process. To mitigate this, for all methods, we independently and randomly repeated the “selection evaluation” process times and reported the averaged results.
V Results and Discussion
V-A Comparison of Accuracy
| Method | GPT-4o-mini | Qwen2.5-Coder-32B | Llama-3.1-70B | DeepSeek-V3 |
| Sampling | 35.25 | 44.00 | 32.00 | 69.00 |
| Sampling+Filtering* | 37.50 | 44.50 | 31.00 | 68.25 |
| Self-Repair* | 33.00 | 42.00 | 29.25 | 60.75 |
| Reflexion* | 37.50 | 48.75 | 25.00 | 62.50 |
| INTERVENOR* | 27.75 | 20.25 | 20.25 | 47.50 |
| CodeCoT | 31.25 | 36.50 | 19.50 | 63.50 |
| AgentCoder | 31.25 | 41.75 | 22.00 | 60.25 |
| MBR-Exec | 33.75 | 45.00 | 38.75 | 70.00 |
| CodeT | 46.25 | 47.50 | 41.25 | 72.50 |
| CoCoEvo | 49.75 | 55.75 | 45.00 | 76.25 |
Denotes that the methods were originally designed to operate with pre-defined test cases. For this work, we adapted these methods to utilize test cases generated by the LLMs.
In this section, we present the results of the accuracy comparison for the methods. Table III reports the pass@1 metric for each method across different LLMs on the LeetCode-Contest dataset. As shown in Table III, our proposed CoCoEvo method achieves the highest performance across all LLMs, demonstrating its superior ability to generate accurate code. This success can be attributed to the co-evolution framework employed by CoCoEvo, which facilitates a more effective iterative refinement between programs and test cases.
An interesting observation concerns the repair-based methods such as Self-Repair, Reflexion, and INTERVENOR. When these methods utilize test cases generated by LLMs instead of pre-defined ones, their effectiveness diminishes significantly. In some instances, their performance is even worse than that of the basic Sampling approach. This is because the erroneous test cases generated by LLMs can mislead these methods by providing inaccurate feedback, causing the repair process to make suboptimal adjustments to the code. Consequently, this results in lower accuracy and poorer overall performance. Similarly, for methods like Sampling+Filtering, where test cases are used to filter and select the best code, performance remains comparable to the Sampling method when the test cases are generated by LLMs. This highlights a critical limitation: these methods struggle to adapt effectively in scenarios where reliable pre-defined test cases are unavailable.
Furthermore, methods that leverage the agreement between programs and test cases, such as MBR-Exec and CodeT, exhibit higher performance compared to Sampling and repair-based approaches. Among these, CodeT stands out, delivering consistent improvements in pass@1 across all LLMs. This suggests that considering the alignment and mutual validation of programs and test cases is a promising strategy for enhancing code generation accuracy.
Fig. 9 illustrates the changes of pass@1 during the evolution process. While fluctuations in performance are observed, the method demonstrates consistently strong performance on average. The figure shows that although the CoCoEvo method starts with a relatively low pass@1, its performance steadily improves over the course of evolution iterations, eventually surpassing that of the CodeT method. This demonstrates the effectiveness of the CoCoEvo method. By alternating the evolution of programs and test cases, the program population progressively converges toward the correct solution.
Fig. 10 illustrates the changes in the accuracy of the test population throughout the evolution process. At initialization, the CoCoEvo method achieves a slightly higher test accuracy than the randomly generated method. Specifically, the CoCoEvo method generates test cases once and extracts test cases, whereas the random method involves generating test cases times and extracting test cases each time. This indicates that as the number of test case generations by LLMs increases, test accuracy tends to decrease slightly.
Interestingly, For the CoCoEvo method, the peak test accuracy is observed in the second iteration. In subsequent iterations, the accuracy declines slightly, potentially due to the Pareto selection process, where test cases with higher discrimination but lower confidence are accepted. This may inadvertently introduce some incorrect test cases. Nevertheless, the decline is gradual, and the test accuracy remains consistently high, making this behavior acceptable.
V-B Comparison of Program Evolution Scheduler
| Crossover Rate Scheduler | Pass@1 |
| Constant | 51.00 |
| Cosine | 55.75 |
To assess the effectiveness of the crossover rate scheduler, we compare the cosine scheduler with a constant scheduler in our experiments, using a mutation rate of and a crossover rate of . These experiments were performed on the Qwen2.5-Coder-32B and the LeetCode-Contest dataset. The results, summarized in Table IV and Figure 11, highlight key differences between the cosine scheduler and the constant rate approach.
At the early stages of the evolution process, the constant method, characterized by a fixed crossover and mutation rate, demonstrates faster convergence and initially outperforms the cosine scheduler due to its higher crossover rate. However, this rapid convergence causes the Constant method to stagnate and become trapped in a local optimum, ultimately limiting its overall performance.
In contrast, the cosine scheduler, despite its lower initial performance, maintains consistent improvement throughout the evolution process. By adapting the crossover rate dynamically, the scheduler avoids premature convergence and achieves superior results in later stages. Notably, it surpasses the performance of the Constant method after the fifth iteration, ultimately delivering a higher pass@1 rate.
This comparison underscores the importance of dynamic scheduling in optimizing program evolution, particularly in scenarios prone to local optima.
V-C Comparison of Program Fitness Function
| Program Fitness Function | Pass@1 |
| Pass Rate | 53.50 |
| CodeT Score | 55.75 |
To evaluate the impact of different program fitness functions, we conducted a comparative analysis using the CodeT score and the simple pass rate as fitness functions. The base model for these experiments was Qwen2.5-Coder-32B with the LeetCode-Contest dataset serving as the benchmark. The results are summarized in Table V and illustrated in Figure 12.
The findings reveal the insight that calculating the pass rate on test cases directly as the fitness function yields a lower pass@1 value compared to utilizing the CodeT score. This behavior can be attributed to the advantages of the CodeT score, which incorporates the congruence between programs and test cases more effectively. By accounting for this alignment, the CodeT score mitigates the negative effects of erroneous or poorly designed test cases on program fitness, leading to more stable and improved performance. These results underscore the superiority of the CodeT score as a program fitness function, especially in scenarios where robustness and convergence are critical.
V-D Comparison of Test Fitness Function
| Test Fitness Function | Pass@1 |
| Failure Rate | 37.00 |
| Pass Rate | 51.50 |
| Confidence | 51.50 |
| Pareto | 55.75 |
In this section, we present the results of comparing various test fitness functions. Specifically, we evaluate Failure Rate, Pass Rate, Confidence, and Pareto as methods for test fitness calculation. The experiments are conducted using the Qwen2.5-Coder-32B model on the LeetCode-Contest dataset. The detailed results are depicted in Fig. 13, and additional information on test case accuracy is provided in Fig. 14.
Among the test fitness functions, Pass Rate directly calculates the proportion of test cases successfully passed by a program, while Confidence represents a weighted pass rate, as described in Section III-C. Similarly, Failure Rate measures the failure rate of a program on the test cases. For these three fitness calculation methods, greedy selection is employed as the offspring selection method.
Our findings indicate that employing the Pareto method as the selection mechanism yields the best performance. The Pareto approach prioritizes test cases with both high discrimination and high confidence levels, thereby demonstrating superior effectiveness. In contrast, using Failure Rate as the fitness function results in a sharp decline in pass@1 performance. Previous studies on co-evolution have generally argued that the confidence level of a test case increases as more programs fail on it. However, we discovered that if the test cases contain errors, this method can lead to a population dominated by faulty test cases, which fail to provide accurate feedback to the programs.
Furthermore, Fig. 13 reveals that employing Confidence as the test fitness function outperforms using Pass Rate by a small margin. The Confidence function weights the fitness of test cases based on program performance, with programs demonstrating higher fitness values contributing more heavily to the confidence score. This approach ensures that more reliable programs have a greater influence on test case evaluation.
V-E Ablation Study
| Method | Pass@1 |
| Sampling | 44.00 |
| w/o test evolution | 51.75 |
| CoCoEvo | 55.75 |
We conducted ablation experiments using the Qwen2.5-Coder-32B model on the LeetCode-Contest dataset to evaluate the effectiveness of the test case evolution module. In these experiments, we removed the test case evolution module and directly utilized the test cases generated by the LLMs to evaluate the code. The comparison results are presented in Table VII and Fig. 15.
The experimental results demonstrate that using only the program evolution module, without incorporating the test case evolution module, still outperforms the baseline sample-based method. However, after integrating the test case evolution module, the performance of the CoCoEvo method is further enhanced. This highlights the critical role of the test case evolution module in the co-evolution process.
Through the co-evolution of test cases, two key benefits are achieved. First, the accuracy of test cases within the population improves, as they are refined based on the agreement among programs. Second, by applying the Pareto method, discriminative and challenging test cases are preserved. These refined test cases provide more accurate and meaningful feedback to the program evolution process, ultimately improving the program’s pass rate in the final results.
VI Conclusion
This paper introduced CoCoEvo, an LLM-based co-evolution framework that advances automated code generation by integrating simultaneous program and test case evolution. By eliminating the reliance on pre-defined test cases, CoCoEvo addresses a critical limitation in existing approaches, particularly in scenarios lacking comprehensive testing resources. The dynamic crossover rate scheduling mechanism and the multi-objective optimization for test case selection were key to the framework’s success, enabling effective exploration and refinement throughout the evolution process. Experimental results on the LeetCode-Contest dataset, utilizing four leading LLMs, demonstrated that CoCoEvo achieved superior performance compared to traditional methods.
Beyond its practical implications in software development, CoCoEvo opens new avenues for integrating evolutionary algorithms with LLMs. Future work will focus on enhancing the scalability of CoCoEvo for project-level code generation and incorporating additional quality metrics for test case evaluation. Furthermore, exploring its application in other domains, such as automated debugging and software maintenance, could further expand its impact.

### v2

@@
 CoCoEvo eliminates the dependency on pre-defined test cases by generating both programs and test cases directly from natural language problem descriptions and function headers.
-The framework employs specialized evolutionary operators, including LLM-based crossover and mutation operators for program evolution, along with a test case generation operator for test case evolution.
+The framework employs specialized evolutionary operators, including LLM-based crossover and mutation operators for program evolution, along with an additional test case generation operator for test case evolution.
 Additionally, we propose optimization strategies such as a crossover rate scheduler to balance exploration and convergence, and a multi-objective optimization method for test case selection.
@@
 Existing approaches frequently rely on pre-defined test cases for evaluation.
-For instance, Sampling+Filtering [5] first generates a batch of candidate codes, evaluates them using pre-defined test cases, and selects the code with the highest score as the final solution.
+For instance, Sampling+Filtering [5] first generates a batch of candidate codes, evaluates them using pre-defined test cases, and selects the code with the highest score as the final answer.
 Other methods, such as Reflexion [6], Self-Repair [7], INTERVENOR [8], LDB [9], and PairCoder [10], utilize execution feedback on pre-defined test cases to refine or repair generated code.
@@
 However, pre-defined test cases are not always available.
-In test-driven development (TDD) [11], test cases are defined before program implementation.
-In real-world scenarios, software development frequently relies on natural language requirements, with no pre-existing test cases or example programs.
+In test-driven development (TDD) [11], test cases need to be defined before program implementation.
+In real-world software development, teams often rely solely on natural language requirements, as there are no pre-existing test cases or example programs.
 Furthermore, creating pre-defined test cases can be labor-intensive, requiring significant manual effort.
@@
 However, studies such as [12, 17] highlight that LLM-generated test cases can contain significant errors, potentially leading to unreliable evaluations and incorrect feedback.
-Moreover, many of these methods neglect to verify the quality of the generated test cases.
-While approaches like CodeT and MBR-Exec implement basic filtering mechanisms for test cases, other methods often use these generated test cases directly for code evaluation without ensuring their correctness.
+Moreover, many of these methods fail to verify the quality of the generated test cases.
+While approaches implement basic filtering mechanisms for test cases, other methods often use generated test cases directly for code evaluation without ensuring their correctness.
 There are also previous co-evolution methods based on code and test cases, which have primarily focused on code maintenance and program repair.
@@
 1).
-In CoCoEvo, both test cases and programs are generated by LLMs, requiring only natural language problem descriptions and function headers.
+In CoCoEvo, both programs and test cases are generated by LLMs based on natural language problem descriptions.
 No pre-defined programs, test cases, or templates are needed.
@@
 In program evolution, we introduce LLM-based crossover and mutation operators to generate program offspring.
-The current population of test cases is used to evaluate program fitness.
+The current test case population is used to calculate program fitness.
 In test case evolution, we design an additional test case generation operator to produce test case offspring.
-Conversely, the current population of programs is used to evaluate test case fitness.
+Conversely, the current program population is used to calculate test case fitness.
 For programs, we implement a fitness calculation method inspired by CodeT [12], which assesses agreement between programs and test cases.
@@
 We conducted experiments on four state-of-the-art models: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4].
-We compared with both basic code generation methods such as Sampling [5] and Sampling+Filtering [5], repair-based methods, and agreement-based methods.
+We compared with both basic code generation methods, repair-based methods, and agreement-based methods.
 The results demonstrate that CoCoEvo outperforms all existing methods.
@@
 For program evolution, we design a dynamic crossover rate scheduler that enhances exploration early on and accelerates convergence later.
-For test cases, we develop a multi-objective optimization method for fitness evaluation and selection.
+For test cases, we develop a multi-objective optimization method for fitness calculation and new population selection.
 3.
@@
 The experimental results demonstrate that CoCoEvo achieves state-of-the-art performance, surpassing existing methods.
-Additionally, we conduct ablation studies to evaluate the utility of each module within the framework, further highlighting the effectiveness of our approach.
+Additionally, we conduct comparative experiments to evaluate the utility of each module within the framework, further highlighting the effectiveness of our approach.
 The remainder of this article is organized as follows.
@@
 Sections IV and V present the experimental settings and discussions, respectively.
-Finally, Section VI concludes this article and outlines future work.
+Finally, Section VI concludes this paper and discusses the limitations and future work.
 II Related Work
 II-A Enhancing Code Generation Accuracy with Test Case
-The integration of test cases has emerged as a pivotal strategy for improving the quality of code generated by LLMs.
-Numerous studies leverage test cases to filter, evaluate, and refine generated programs, leading to higher-performing and more reliable code.
-Below, we provide an overview of notable methodologies in this domain.
+Integrating test cases is now crucial for improving LLM-generated code. Key methodologies use them to filter, evaluate, and refine programs, enhancing performance and reliability.
 Li et al.
-[5] introduce a Sampling+Filtering approach, where a large set of candidate programs is sampled and scored based on performance against a provided test set.
+[5]’s Sampling+Filtering approach samples numerous candidate programs and scores them against a given test set.
 Chen et al.
-[12] propose a framework in which both code and test cases are generated by LLMs simultaneously, using a dual agreement metric to identify and select better programs.
+[12] use LLMs to co-generate code and tests, selecting superior programs via a dual agreement metric.
 Shi et al.
-[13] assess semantic similarity by executing generated programs on test cases and leveraging the execution results to guide program selection.
-Beyond evaluation, test cases play a critical role in iterative repair and refinement of generated code.
-Execution outcomes from test cases can be used to repair faulty programs.
-For instance, Shinn et al.
-[6] introduce the Reflexion framework, where failure messages from test executions are converted into verbal feedback for program repair.
+[13] use test execution results to assess program similarity, guiding program selection.
+Beyond evaluation, test-case results drive iterative repair and refinement of generated code.
+Reflexion [6] converts test failures into verbal feedback for program repair.
 Similarly, Olausson et al.
@@
 [9] extend this idea by enabling step-by-step debugging.
-Huang et al.
-[14] adopt a chain-of-thought prompting [15] approach that generates code and tests sequentially, iteratively improving the code based on test feedback.
+CodeCoT [14] uses chain-of-thought prompting [15] to sequentially generate code and tests, iteratively refining the code with test feedback.
 Zhang et al.
 [10] introduce a pair programming framework where test feedback informs repair strategies and planning.
-The generation of test cases differs across these methodologies.
-For example, CodeT [12] and CodeCoT [14] rely entirely on LLMs to generate test cases, whereas Reflexion employs LLMs primarily for simple programming tasks.
-Most other methods, however, assume pre-defined test cases as part of their evaluation frameworks, leveraging them for filtering, refinement, and program selection.
 II-B Automated Test Case Generation
 Early works, such as [24], leveraged search-based methods to generate test cases.
-More recently, there has been a growing trend toward utilizing LLMs for test case generation.
+More recently, research has shifted toward using LLMs for this task.
 Studies such as [25], [26], and [27] have explored the effectiveness of LLMs in automating program testing.
 Similarly, [28] investigates methods for enhancing existing test cases using LLMs.
-Other approaches, such as TOGA [29], TOGLL [30], and the method described in [31], focus on generating test oracles based on predefined templates.
-Meanwhile, [32] introduces a model trained with deep reinforcement learning for text-to-test case generation.
+Others, like TOGA [29], TOGLL [30], and [31], generate test oracles from predefined templates.
+Meanwhile, [32] trained a deep reinforcement learning model for text-to-test case generation.
 The application of LLMs in software development extends to generating test cases for various scenarios.
@@
 Similarly, [34] demonstrates the use of LLMs to create test cases aimed at reproducing general software bugs.
-LLMs are also integrated into code generation studies to improve the accuracy of generated code through test case generation.
+LLMs are also employed to improve the code accuracy through test case generation.
 CodeT [12], for example, utilizes LLMs with zero-shot prompts to directly generate test cases.
@@
 The integration of LLMs with evolutionary algorithms has been extensively explored to enhance code generation.
-For example, Bradley et al.
+Bradley et al.
 [35] introduced OpenELM, an open-source Python library that designs specialized evolutionary operators for code generation.
 Chen et al.
-[36] utilized LLMs as adaptive mutation and crossover operators in evolutionary neural architecture search (NAS), optimizing neural network design by merging LLM capabilities with evolutionary processes.
+[36] employed LLMs as adaptive operators for evolutionary neural architecture search to optimize network designs.
 Liu et al.
-[37] proposed an automated approach for evolving optimization algorithms using LLMs, which eliminates the need for manual design by applying standard evolutionary operations such as initialization, selection, and mutation.
-In a related study, Liu et al.
-[38] developed a framework for evolving heuristics by integrating LLMs with evolutionary computation to co-evolve both natural language descriptions and executable code for heuristic design, further extending previous methodologies.
+[37] proposed using LLMs to evolve optimization algorithms, automating the processes of initialization, selection, and mutation.
+Liu et al.
+[38] introduced a framework integrating LLMs with evolutionary computation to co-evolve heuristic descriptions and executable code.
 Ma et al.
-[39] introduced an algorithm combining LLMs with evolutionary search to generate and optimize reward functions for reinforcement learning tasks, incorporating a reward reflection mechanism for refinement.
+[39] combined LLMs with evolutionary search to generate and refine RL reward functions via a reward-reflection mechanism.
 Meyerson et al.
-[40] demonstrated the use of few-shot prompting with LLMs to generate offspring in evolutionary algorithms, enabling a general crossover operator by concatenating parent solutions into a single prompt, which can adapt to various domains.
+[40] used few-shot prompting to create a domain-agnostic crossover operator by concatenating parent solutions into one prompt.
 Similarly, Romera et al.
-[41] integrated LLMs with evolutionary algorithms to create context-aware program mutations, eliminating the need for manually defined mutation operators.
+[41] integrated LLMs with evolutionary algorithms to create context-aware program mutations.
 Ye et al.
-[42] proposed ReEvo, which combines evolutionary search with LLM-driven reflections to generate optimization heuristics, efficiently exploring heuristic spaces while offering verbal gradients for optimization.
-Lastly, AutoTest [43] combines automated test case generation with code execution in an evolutionary framework, utilizing LLMs to co-generate code solutions and their corresponding test cases for enhanced optimization.
+[42]’s ReEvo combines evolutionary search and LLM-driven reflections to generate heuristics via verbal optimization gradients.
+AutoTest [43] embeds LLM co-generation of code and test cases in an evolutionary loop for improved optimization.
 Additionally, various LLM-based co-evolution methods have been proposed.
@@
 Chi et al.
-[19] leveraged LLMs with dynamic validation to automate co-evolution between production and test code, identifying obsolete test cases to improve software quality and reduce maintenance overhead.
-TestART [44] enhances unit testing for LLMs by combining automated test generation with repair, using template-based fixes and coverage feedback to improve test accuracy.
+[19] used LLMs with dynamic validation to automate production and test code co-evolution, identifying obsolete tests to enhance software quality and lower maintenance overhead.
+TestART [44] enhances LLM unit testing by automatically generating and repairing tests, using template-based fixes and coverage feedback to improve accuracy.
 III Methods
@@
 2.
-The evolutionary process begins with the random initialization of both the program population and the test case population, achieved through the random generating programs and test cases using LLMs.
+The process begins by randomly initializing both the program and test-case populations, using LLMs to generate the initial programs and test cases.
 The co-evolution process follows a loop consisting of two alternating steps: program evolution and test case evolution, each of which includes a Generation-Evaluation-Selection sequence.
@@
 Conversely, during test case evolution, the program population is utilized to evaluate test cases.
-For programs, we introduce LLM-based crossover and LLM-based mutation operators for offspring generation.
+For programs, we introduce LLM-based crossover and mutation operators for offspring generation.
 For test cases, we implement an LLM-based additional test case generation operator.
-Once the preset number of iterations is reached, the co-evolution loop terminates, and the program with the highest fitness score is returned as the final solution.
+Once the preset number of iterations is reached, the co-evolution loop terminates, and the program with the highest fitness score is returned as the final answer.
 Algorithm 1 provides the pseudo-code detailing the CoCoEvo process.
@@
 III-B Program Evolution
-This section provides a detailed description of strategies employed in program evolution.
-Program fitness function. During evaluation, programs and test cases undergo cross-evaluation, as illustrated in Fig. 3. The outcomes are represented by a matrix :
-We compute the confidence score of each program, adopting the approach in [12].
+Program evaluation. During evaluation, programs and test cases are cross-evaluated, as illustrated in Fig. 3. The outcomes are represented by a matrix :
+where and represent the ‑th program and the ‑th test case, respectively.
+We design a Confidence metric to evaluate program fitness, drawing inspiration from CodeT [12].
 Programs that pass the same set of test cases are grouped into a set , while the corresponding test cases are grouped into a set .
-The confidence of program is then defined as
-where represents the set containing program , and represents the associated test set. The confidence score evaluates the likelihood of a program being correct. The fitness of program is defined as
+The Confidence of program is defined as
+where the size of and are denoted by and , respectively. This Confidence score evaluates the likelihood of a program being correct. The fitness of program is defined as
 where a higher fitness score indicates a higher probability of correctness for the program.
-Crossover Rate Scheduler To address the multimodal nature of the program search space and mitigate the risk of premature convergence, a dynamic scheduler is employed to regulate crossover and mutation rates.
-A higher mutation rate is applied in the early stages to promote exploration, while the mutation rate decreases, and the crossover rate increases in later stages to refine convergence.
+Crossover Rate Scheduler.
+When leveraging LLMs to perform program crossover and mutation, the crossover operation typically synthesizes the strengths of disparate programs but may drive the population toward premature convergence, whereas the mutation operation often introduces novel features yet cannot guarantee evolution in a more desirable direction.
+Therefore, we propose a dynamic crossover rate scheduler that adaptively adjusts the number of crossover and mutation operations.
+We define the crossover rate as the proportion of offspring programs generated through the crossover operator.
+In the early stages, a lower crossover rate is applied to promote exploration, and the crossover rate increases in later stages to refine convergence.
+This enables extensive exploration of the search space during the early stages of evolution, while facilitating rapid convergence toward superior solutions in later generations.
 Unlike traditional random mutation methods, the LLM-based program mutation operator leverages LLMs for code generation, which could retain key features of high-quality programs.
 This allows for a higher mutation rate without significant loss of desirable traits.
-We adopt the cosine annealing method [45] for scheduling crossover rate:
-This formulation ensures that the crossover rate increases smoothly from to . The number of crossover operations and the number of mutation operations are defined as follows:
-where is the program population size, and represent the number of program crossover and mutation operations in iteration round , respectively.
-When the initialization is completed, the iteration round is set to .
-After initialization, the iteration round starts at and increments by after a co-evolution loop.
+Thus, we adopt the cosine annealing method [45] for scheduling crossover rate:
+where denotes the crossover rate at the -th iteration, and represent the initial crossover rate and the final crossover rate, which are set to and , respectively.
+During the iteration process, starts at and ends at .
+In the formula, both and are decremented by , since in the first iteration () the program population is doing initialization and no crossover or mutation operations are performed.
+This formulation ensures that the crossover rate increases smoothly from to .
+In each generation, offspring programs are produced.
+The number of crossover operations is determined by multiplying the crossover rate by and taking the floor of the result.
+The number of mutation operations is equal to minus the number of crossover operations, as calculated by the following equations:
+where is the program population size, and represent the number of program crossover and mutation operations in the -th iteration, respectively.
 Fig.
-4 illustrates the variation in crossover rates over evolutionary rounds.
+4 presents an example of the variation in crossover rate over iterations.
+Since when , the population is randomly initialized, the figure illustrates changes in the crossover rate from to , with the parameter set to .
 This approach ensures a smooth transition between exploration (early stages) and convergence (later stages), enabling effective optimization of the solution space.
 III-C Test Case Evolution
-Test fitness function. The evaluation of test cases involves two key metrics: confidence and discrimination. The confidence of test case is calculated as
-where represents the total number of programs, represents the fitness of program , and is the evaluation matrix indicating whether passed .
+Test case evaluation.
+The evaluation of test cases involves two key metrics: Confidence and Discrimination.
+Confidence evaluates the probability that a test case is correct, and discriminative power evaluates a test case’s ability to distinguish between programs.
+The Confidence of test case is calculated as
+where represents the size of the program population, represents the fitness of program , and is the evaluation matrix indicating whether passed .
 Intuitively, Confidence measures the degree of agreement between the program population and the test cases, with higher program fitness contributing more weight to the degree of agreement.
@@
 This indicator assesses whether the test cases can exert evolutionary pressure on the program population.
-The selection of test cases is performed using the Pareto approach for multi-objective optimization [22], as illustrated in Fig.
+Test case selection.
+The selection of test cases is performed using the Pareto approach for multi-objective optimization [22].
+The Pareto frontier, also known as the Pareto front or Pareto boundary, represents the set of non-dominated solutions in a multi-objective optimization problem.
+A solution is considered Pareto optimal if no other feasible solution improves one objective without causing a deterioration in at least one other.
+Formally, given a vector-valued objective function , a solution is Pareto optimal if there does not exist another solution such that for all , with at least one strict inequality.
+In test case selection, the Pareto front is constructed based on the two metrics, and , which represent the Confidence of test cases and the Discrimination of test cases.
+The Pareto front is illustrated in Fig.
 5.
-The Pareto front is constructed based on the two optimization metrics, confidence and discrimination.
-Test cases with high confidence and low discrimination represent basic test cases that cover common scenarios, with a high probability of being correct.
-Test cases with high discrimination represent more challenging cases that are effective at exposing errors.
-We noticed that there are some test cases with very low confidence levels that sometimes have a very high degree of discrimination.
-To avoid this type of test case rom being selected, the test cases with confidence lower than the average in the selection results will be filtered out.
+Test cases with high and low represent basic test cases that cover common scenarios, with a high probability of being correct.
+Test cases with high represent more challenging cases that are effective at exposing errors.
+During the selection process, preference is given to individuals that perform well on both indicators.
+This strategy aims to: (1) preserve the correctness of the test case population, ensuring accurate evaluation of the program, and (2) retain test cases with high discriminative power to maintain selection pressure on the program population.
+We also noticed that there are some test cases with very low that sometimes have a relatively high .
+To avoid adverse effects of these test cases on program evolution, the test cases with lower than the average in the selection results will be filtered out.
 III-D LLM-based Evolution Operators
-Program crossover operator.
-We utilize LLMs to perform program crossover, as illustrated in Fig. 6. The process involves the following steps:
+Program crossover operator. We utilize LLMs to perform program crossover, as illustrated in Fig. 6. The process involves the following steps:
 1.
@@
 Information Construct.
-The prompt contains the following information: The entire existing population of test cases, the best program the program population and the line coverage information on it, where covered lines are marked as “[+]” and uncovered lines as “[-]”.
+The prompt contains the following information: The entire existing population of test cases, the best program in the program population, and the line coverage information on it, where covered lines are marked as “[+]” and uncovered lines as “[-]”.
 2.
@@
 We collected a dataset of programming problems from the weekly and biweekly contests hosted on the LeetCode platform 111https://leetcode.com/, forming the LeetCode-Contest dataset.
-In the dataset, each problem includes a prompt comprising the function header and a natural language description in the form of a docstring, along with test cases and reference solutions.
+In the dataset, each problem includes a prompt comprising the function header and a natural language description in the form of a docstring, along with program solutions and approximately ground truth test cases per problem.
+These comprehensive ground truth test cases can closely simulate the real-world scenario of submissions on LeetCode, making our evaluation more convincing.
 To ensure data integrity and avoid potential leakage, all selected problems were released after March 2024.
@@
 Specifically, we removed all test input-output samples from the problem descriptions, retaining only the natural language problem statements, the function header, and the associated constraints.
-For each problem, we collected Python solution codes and a substantial set of test cases available on the website. The detailed information of the LeetCode-Contest dataset is summarized in Table I.
-| Problems | 80 |
-| Average Test Cases | 644.40 |
-| Selected Contests | weekly-contest-402, weekly-contest-401, biweekly-contest-132, weekly-contest-400, weekly-contest-399, biweekly-contest-131, weekly-contest-398, weekly-contest-397, biweekly-contest-130, weekly-contest-396, weekly-contest-395, biweekly-contest-129, weekly-contest-394, weekly-contest-393, biweekly-contest-128, weekly-contest-392, weekly-contest-391, biweekly-contest-127, weekly-contest-390, weekly-contest-389 |
+The detailed information of the LeetCode-Contest dataset is summarized in Table I.
+Problems | 80 | ||||||||||
+Average Test Cases | 644.40 | ||||||||||
+Selected Contests |
 IV-B Models
-We utilized four widely adopted and powerful models for our experiments: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4].
+We utilized four widely adopted and powerful LLMs for our experiments: GPT-4o-mini [1], Qwen2.5-Coder-32B [3], Llama-3.1-70B [2], and DeepSeek-V3 [4].
 For GPT-4o-mini, we accessed the API via the OpenAI platform 222https://platform.openai.com/.
@@
 1.
-Sampling [5]. A basic method that generates a large number of code samples and randomly selects one as the final answer.
+Sampling. A basic method that generates a large number of code samples and randomly selects one as the final answer.
 2.
@@
 The clustering score is calculated as the product of the number of codes in a cluster and the number of test cases passed.
-For all baselines, if the method requires test cases for evaluation or feedback generation, the test cases are generated using the same LLMs that perform code generation.
+For the baselines (Sampling+Filtering, Self-Repair, Reflexion, INTERVENOR) that originally relied on pre-defined test cases for program evaluation, we adapt them to our setting by replacing the pre-defined test cases with LLM self-generated test cases.
 IV-D Performance Indicators
-The ground truth test cases in the dataset are used to evaluate the accuracy of the generated code.
-A code is considered correct if it passes all the test cases.
+The ground truth test cases in the dataset are used to evaluate the accuracy of the generated program.
+A code is considered correct if it passes all the ground truth test cases.
 The performance of the approaches on the dataset is measured using the pass@1 metric [46].
@@
 IV-E Experimental Settings
-Given that invoking LLMs for code generation or repair consumes significantly more time and resources compared to executing code, we evaluated the performance of all methods under the condition that the number of LLM calls for code generation or repair is held constant.
-To ensure practical usability and align with the parameter settings of most code generation methods, we set the number of LLM calls for code generation to iterations.
-For methods requiring LLM-generated test cases, we limited test case generation to iterations, with a maximum of test cases extracted per iteration.
-The parameter settings of the CoCoEvo method can be found in Table II.
-| Program Population Size | 10 |
-| Iterations | 10 |
-| Min Test Case Population Size | 10 |
+Following the comparison settings used in prior studies [7, 13, 12], we compare the performance of different methods by constraining the number of code generation attempts made by the LLMs to be equal.
+Based on the common limitations of these approaches and taking into account the cost of invoking LLMs, we limit the number of LLM code generation calls to .
+For methods that require test case generation, we impose a similar constraint by limiting the number of test case generation calls made to the LLMs to , with a maximum of test cases extracted per generation.
+For CoCoEvo, we adopt a balanced configuration by setting both the population size and the number of iterations to .
+Comparative experiments with alternative population sizes and detailed parameter settings of baselines and CoCoEvo can be found in the supplementary material.
+Since different code generation strategies may require varying numbers of tokens, we provide a detailed analysis of token consumption in Section V-G.
 We observed that when calculating the pass@1 metric, if multiple codes achieve the same score, one code will be randomly selected as the final result.
@@
 In this section, we present the results of the accuracy comparison for the methods.
-Table III reports the pass@1 metric for each method across different LLMs on the LeetCode-Contest dataset.
-As shown in Table III, our proposed CoCoEvo method achieves the highest performance across all LLMs, demonstrating its superior ability to generate accurate code.
+Table II reports the pass@1 metric for each method across different LLMs on the LeetCode-Contest dataset.
+As shown in Table II, our proposed CoCoEvo method achieves the highest performance across all LLMs, demonstrating its superior ability to generate accurate code.
 This success can be attributed to the co-evolution framework employed by CoCoEvo, which facilitates a more effective iterative refinement between programs and test cases.
-An interesting observation concerns the repair-based methods such as Self-Repair, Reflexion, and INTERVENOR.
+An interesting observation concerns the code repair based methods such as Self-Repair, Reflexion, and INTERVENOR.
 When these methods utilize test cases generated by LLMs instead of pre-defined ones, their effectiveness diminishes significantly.
 In some instances, their performance is even worse than that of the basic Sampling approach.
-This is because the erroneous test cases generated by LLMs can mislead these methods by providing inaccurate feedback, causing the repair process to make suboptimal adjustments to the code.
-Consequently, this results in lower accuracy and poorer overall performance.
-Similarly, for methods like Sampling+Filtering, where test cases are used to filter and select the best code, performance remains comparable to the Sampling method when the test cases are generated by LLMs.
-This highlights a critical limitation: these methods struggle to adapt effectively in scenarios where reliable pre-defined test cases are unavailable.
-Furthermore, methods that leverage the agreement between programs and test cases, such as MBR-Exec and CodeT, exhibit higher performance compared to Sampling and repair-based approaches.
-Among these, CodeT stands out, delivering consistent improvements in pass@1 across all LLMs.
-This suggests that considering the alignment and mutual validation of programs and test cases is a promising strategy for enhancing code generation accuracy.
+This is primarily due to the faulty LLM-generated tests mislead the repair process, leading to suboptimal code modifications and degraded overall performance.
+Similarly, Sampling+Filtering yields only a slight improvement over plain Sampling under these conditions, highlighting these approaches’ dependence on reliable test suites.
+Furthermore, methods that leverage the agreement between programs and test cases, such as MBR-Exec and CodeT, outperform both Sampling and repair-based methods.
+Among these, CodeT stands out, delivering consistent improvements in pass@1 across all LLMs, suggesting that mutual validation of code and test cases is a promising strategy for improving generation accuracy.
 Fig.
@@
 Fig.
-10 illustrates the changes in the accuracy of the test population throughout the evolution process.
-At initialization, the CoCoEvo method achieves a slightly higher test accuracy than the randomly generated method.
-Specifically, the CoCoEvo method generates test cases once and extracts test cases, whereas the random method involves generating test cases times and extracting test cases each time.
-This indicates that as the number of test case generations by LLMs increases, test accuracy tends to decrease slightly.
-Interestingly, For the CoCoEvo method, the peak test accuracy is observed in the second iteration.
-In subsequent iterations, the accuracy declines slightly, potentially due to the Pareto selection process, where test cases with higher discrimination but lower confidence are accepted.
-This may inadvertently introduce some incorrect test cases.
-Nevertheless, the decline is gradual, and the test accuracy remains consistently high, making this behavior acceptable.
+10 presents the accuracy trends of the test case population and offspring over the evolutionary iterations, as well as the accuracy of the randomly generated test cases.
+As shown in the figure, the accuracy of the test case population consistently exceeds that of randomly generated test cases, demonstrating the effectiveness of the alternating selection strategy in the co-evolutionary framework.
+This indicates that selecting test cases based on their Confidence is beneficial for identifying faulty test cases.
+Moreover, the strategy for generating additional test cases exhibits a modest upward trend in the accuracy of offspring test cases, which also contributes positively to maintaining the overall accuracy of the test case population.
+Interestingly, the peak test accuracy of the test population is observed in the second iteration.
+In subsequent iterations, the accuracy declines slightly.
+The slight decline can be attributed to the presence of a few incorrect test cases on the Pareto front that exhibit high discriminative capability but low Confidence, thus affecting overall accuracy.
+Nevertheless, the Pareto selection approach preserves a substantial number of accurate and highly discriminative test cases, thereby offering greater benefits to the co-evolutionary process.
+Consequently, the marginal reduction in accuracy remains within an acceptable range.
 V-B Comparison of Program Evolution Scheduler
-| Crossover Rate Scheduler | Pass@1 |
-| Constant | 51.00 |
-| Cosine | 55.75 |
+Model | Crossover Rate | Pass@1 |
+GPT-4o-mini | Constant | 45.00 |
+Cosine | 49.75 | |
+Qwen2.5-Coder-32B | Constant | 51.00 |
+Cosine | 55.75 | |
+Llama-3.1-70B | Constant | 42.75 |
+Cosine | 45.00 | |
+DeepSeek-V3 | Constant | 72.50 |
+Cosine | 76.25 |
 To assess the effectiveness of the crossover rate scheduler, we compare the cosine scheduler with a constant scheduler in our experiments, using a mutation rate of and a crossover rate of .
-These experiments were performed on the Qwen2.5-Coder-32B and the LeetCode-Contest dataset.
-The results, summarized in Table IV and Figure 11, highlight key differences between the cosine scheduler and the constant rate approach.
-At the early stages of the evolution process, the constant method, characterized by a fixed crossover and mutation rate, demonstrates faster convergence and initially outperforms the cosine scheduler due to its higher crossover rate.
+These experiments were performed on the four LLMs and the LeetCode-Contest dataset.
+The results, summarized in Table III and Figure 11, highlight key differences between the cosine scheduler and the constant rate approach.
+At the early stages of the evolution process, the constant method, characterized by a fixed crossover and mutation rate, demonstrates faster convergence and initially outperforms the cosine scheduler on Qwen2.5-Coder-32B, Llama-3.1-70B, and DeepSeek-V3.
 However, this rapid convergence causes the Constant method to stagnate and become trapped in a local optimum, ultimately limiting its overall performance.
 In contrast, the cosine scheduler, despite its lower initial performance, maintains consistent improvement throughout the evolution process.
+On GPT-4o-mini, the cosine scheduler consistently outperforms the method with a constant crossover rate.
 By adapting the crossover rate dynamically, the scheduler avoids premature convergence and achieves superior results in later stages.
@@
 V-C Comparison of Program Fitness Function
-| Program Fitness Function | Pass@1 |
-| Pass Rate | 53.50 |
-| CodeT Score | 55.75 |
+Model | Program Fitness Function | Pass@1 |
+GPT-4o-mini | Pass Rate | 45.50 |
+CodeT Score | 49.75 | |
+Qwen2.5-Coder-32B | Pass Rate | 53.50 |
+CodeT Score | 55.75 | |
+Llama-3.1-70B | Pass Rate | 41.00 |
+CodeT Score | 45.00 | |
+DeepSeek-V3 | Pass Rate | 73.25 |
+CodeT Score | 76.25 |
 To evaluate the impact of different program fitness functions, we conducted a comparative analysis using the CodeT score and the simple pass rate as fitness functions.
-The base model for these experiments was Qwen2.5-Coder-32B with the LeetCode-Contest dataset serving as the benchmark.
-The results are summarized in Table V and illustrated in Figure 12.
-The findings reveal the insight that calculating the pass rate on test cases directly as the fitness function yields a lower pass@1 value compared to utilizing the CodeT score.
-This behavior can be attributed to the advantages of the CodeT score, which incorporates the congruence between programs and test cases more effectively.
-By accounting for this alignment, the CodeT score mitigates the negative effects of erroneous or poorly designed test cases on program fitness, leading to more stable and improved performance.
-These results underscore the superiority of the CodeT score as a program fitness function, especially in scenarios where robustness and convergence are critical.
+These experiments are performed on the four LLMs and the LeetCode-Contest dataset.
+The results are summarized in Table IV and illustrated in Figure 12.
+The results indicate that using the raw pass rate as the fitness function yields a lower pass@1 than the CodeT score.
+This is because CodeT better captures the alignment between programs and test cases, reducing the impact of flawed or poorly designed tests on fitness.
+Consequently, CodeT delivers more stable performance and faster convergence.
+These findings highlight CodeT’s superiority as a fitness function, particularly when robustness and convergence are essential.
 V-D Comparison of Test Fitness Function
-| Test Fitness Function | Pass@1 |
-| Failure Rate | 37.00 |
-| Pass Rate | 51.50 |
-| Confidence | 51.50 |
-| Pareto | 55.75 |
+Model | Test Fitness Function | Pass@1 |
+GPT-4o-mini | Failure Rate | 37.50 |
+Pass Rate | 47.00 | |
+Confidence | 45.50 | |
+Pareto | 49.75 | |
+Qwen2.5-Coder-32B | Failure Rate | 37.00 |
+Pass Rate | 51.50 | |
+Confidence | 51.50 | |
+Pareto | 55.75 | |
+Llama-3.1-70B | Failure Rate | 26.00 |
+Pass Rate | 37.50 | |
+Confidence | 36.75 | |
+Pareto | 45.00 | |
+DeepSeek-V3 | Failure Rate | 66.75 |
+Pass Rate | 71.75 | |
+Confidence | 72.25 | |
+Pareto | 76.25 |
 In this section, we present the results of comparing various test fitness functions.
 Specifically, we evaluate Failure Rate, Pass Rate, Confidence, and Pareto as methods for test fitness calculation.
-The experiments are conducted using the Qwen2.5-Coder-32B model on the LeetCode-Contest dataset.
+These experiments are performed on the four LLMs and the LeetCode-Contest dataset.
 The detailed results are depicted in Fig.
@@
 For these three fitness calculation methods, greedy selection is employed as the offspring selection method.
-Our findings indicate that employing the Pareto method as the selection mechanism yields the best performance.
-The Pareto approach prioritizes test cases with both high discrimination and high confidence levels, thereby demonstrating superior effectiveness.
+Experimental results indicate that employing the Pareto method as the selection mechanism yields the best performance.
 In contrast, using Failure Rate as the fitness function results in a sharp decline in pass@1 performance.
-Previous studies on co-evolution have generally argued that the confidence level of a test case increases as more programs fail on it.
+Previous studies on co-evolution have generally argued that the strength of a test case increases as more programs fail on it.
 However, we discovered that if the test cases contain errors, this method can lead to a population dominated by faulty test cases, which fail to provide accurate feedback to the programs.
 Furthermore, Fig.
-13 reveals that employing Confidence as the test fitness function outperforms using Pass Rate by a small margin.
-The Confidence function weights the fitness of test cases based on program performance, with programs demonstrating higher fitness values contributing more heavily to the confidence score.
-This approach ensures that more reliable programs have a greater influence on test case evaluation.
-V-E Ablation Study
-| Method | Pass@1 |
-| Sampling | 44.00 |
-| w/o test evolution | 51.75 |
-| CoCoEvo | 55.75 |
-We conducted ablation experiments using the Qwen2.5-Coder-32B model on the LeetCode-Contest dataset to evaluate the effectiveness of the test case evolution module.
+13 reveals that employing Confidence as the test fitness function slightly outperforms using Pass Rate in most cases.
+By weighting test cases according to program performance, higher-performing programs contribute more to the Confidence score, ensuring that the most reliable programs have the greatest influence on evaluation.
+V-E Analysis of Test Coverage Impact
+Method | GPT-4o-mini | Qwen2.5-Coder-32B | Llama-3.1-70B | DeepSeek-V3 |
+Sampling+Filtering | 37.50 | 44.50 | 31.00 | 68.25 |
+Sampling+Filtering w/ cov | 36.25 | 47.75 | 33.25 | 68.50 |
+MBR-Exec | 33.75 | 45.00 | 38.75 | 70.00 |
+MBR-Exec w/ cov | 35.00 | 45.00 | 33.75 | 72.50 |
+CodeT | 46.25 | 47.50 | 41.25 | 72.50 |
+CodeT w/ cov | 47.75 | 50.25 | 39.00 | 71.50 |
+CoCoEvo w/o cov | 48.25 | 53.00 | 42.75 | 76.00 |
+CoCoEvo | 49.75 | 55.75 | 45.00 | 76.25 |
+During offspring test case generation, we incorporate coverage information of the current test case population on the best-performing program individual to guide the generation of more targeted test cases.
+To explore whether this strategy could improve baseline performance, we applied it to several compatible baselines.
+Since this strategy requires an initial set of programs and then iteratively generates test cases based on their coverage, we adopt this method for Sampling, Sampling+Filtering [5], MBR-Exec [13], and CodeT [12].
+We also conducted experiments to evaluate the performance of CoCoEvo without the coverage information.
+The results are shown in Table VI.
+The experiments reveal that simply adding test coverage information to baseline methods can sometimes negatively affect performance, as observed in GPT-4o-mini with Sampling+Filtering, Llama-3.1-70B with MBR-Exec and CodeT, and DeepSeek-V3 with CodeT.
+In contrast, CoCoEvo benefited consistently from coverage information across all LLMs, underscoring that such data is most effective when used in conjunction with the co-evolutionary framework.
+V-F Ablation Study
+Model | Method | Pass@1 |
+GPT-4o-mini | Sampling | 35.25 |
+w/o test evolution | 48.00 | |
+CoCoEvo | 49.75 | |
+Qwen2.5-Coder-32B | Sampling | 44.00 |
+w/o test evolution | 51.75 | |
+CoCoEvo | 55.75 | |
+Llama-3.1-70B | Sampling | 32.00 |
+w/o test evolution | 43.00 | |
+CoCoEvo | 45.00 | |
+DeepSeek-V3 | Sampling | 69.00 |
+w/o test evolution | 74.50 | |
+CoCoEvo | 76.25 |
+We conduct ablation experiments across the four LLMs on the LeetCode-Contest dataset to evaluate the effectiveness of the test case evolution module.
 In these experiments, we removed the test case evolution module and directly utilized the test cases generated by the LLMs to evaluate the code.
@@
 15.
-The experimental results demonstrate that using only the program evolution module, without incorporating the test case evolution module, still outperforms the baseline sample-based method.
-However, after integrating the test case evolution module, the performance of the CoCoEvo method is further enhanced.
+The experimental results demonstrate that using only the program evolution module, without incorporating the test case evolution module, still outperforms the Sampling baseline.
+However, after integrating the test case evolution module, the performance of CoCoEvo is further enhanced.
 This highlights the critical role of the test case evolution module in the co-evolution process.
@@
 These refined test cases provide more accurate and meaningful feedback to the program evolution process, ultimately improving the program’s pass rate in the final results.
+V-G Analysis of Token Usage
+Method | GPT-4o-mini | Qwen2.5-Coder-32B | Llama-3.1-70B | DeepSeek-V3 | |||||||
+Prompt | Completion | Prompt | Completion | Prompt | Completion | Prompt | Completion | ||||
+Sampling | 25823.75 | 14326.96 | 26088.75 | 15082.05 | 25900.00 | 13754.51 | 25821.25 | 19535.79 | |||
+Sampling+Filtering | 28767.25 | 18115.60 | 28932.63 | 18904.11 | 28859.50 | 16305.21 | 28763.38 | 25686.94 | |||
+Self-Repair | 40383.83 | 23086.20 | 40033.05 | 25913.49 | 39661.24 | 23284.41 | 41461.88 | 32701.89 | |||
+Reflexion | 221277.83 | 48270.31 | 236637.13 | 83716.98 | 234549.55 | 54254.50 | 238286.46 | 57093.93 | |||
+INTERVENOR | 107996.59 | 79492.75 | 105999.59 | 93253.63 | 131700.33 | 84377.99 | 102779.86 | 85323.98 | |||
+CodeCoT | 44524.69 | 20815.48 | 49635.12 | 19936.05 | 38194.79 | 22017.59 | 41806.95 | 18614.76 | |||
+AgentCoder | 73915.88 | 35099.35 | 81683.24 | 33976.43 | 71640.85 | 39185.11 | 62136.31 | 27798.86 | |||
+MBR-Exec | 28767.25 | 18115.60 | 28932.63 | 18904.11 | 28859.50 | 16305.21 | 28763.38 | 25686.94 | |||
+CodeT | 28767.25 | 18115.60 | 28932.63 | 18904.11 | 28859.50 | 16305.21 | 28763.38 | 25686.94 | |||
+CoCoEvo Stop-1 | 13558.04 | 5706.66 | 11414.38 | 5404.08 | 15792.08 | 7535.93 | 10031.14 | 5538.18 | |||
+CoCoEvo Stop-2 | 20403.58 | 8255.95 | 17998.85 | 8118.11 | 23626.14 | 10699.68 | 15988.85 | 8192.44 | |||
+CoCoEvo Stop-3 | 26739.18 | 10559.31 | 25055.68 | 11095.44 | 30715.03 | 13666.15 | 22749.20 | 11127.73 | |||
+CoCoEvo Stop-4 | 33405.58 | 12959.24 | 32192.81 | 14077.55 | 38130.99 | 16786.09 | 30061.74 | 14264.30 | |||
+CoCoEvo Stop-5 | 40120.51 | 15358.71 | 39038.18 | 16813.46 | 44900.06 | 19619.78 | 37406.18 | 17349.59 | |||
+CoCoEvo | 61567.61 | 22947.25 | 63638.83 | 26393.06 | 60122.26 | 26128.68 | 66190.20 | 28926.00 |
+Method | GPT-4o-mini | Qwen2.5-Coder-32B | Llama-3.1-70B | DeepSeek-V3 |
+CoCoEvo Stop-1 | 47.75 | 48.75 | 39.75 | 74.75 |
+CoCoEvo Stop-2 | 47.50 | 51.50 | 39.50 | 76.00 |
+CoCoEvo Stop-3 | 46.00 | 51.00 | 42.75 | 75.25 |
+CoCoEvo Stop-4 | 48.00 | 54.00 | 44.00 | 76.00 |
+CoCoEvo Stop-5 | 48.75 | 54.00 | 43.00 | 75.25 |
+CoCoEvo | 49.75 | 55.75 | 45.00 | 76.25 |
+Some strategies, such as reflection and repair employed in some baseline methods, as well as the crossover, mutation, and additional test case generation strategies used in CoCoEvo, may consume more tokens than others.
+We therefore report each method’s token consumption to compare their resource efficiency.
+Table VIII shows the average token consumption per problem on the LeetCode-Contest dataset.
+We report input (Prompt) and output (Completion) tokens separately, as output tokens are generally more costly.
+Notably, we designed an early stopping strategy to investigate whether the token consumption of the CoCoEvo method could be optimized.
+Since problem difficulty varies, simpler problems often require fewer iterations, whereas more complex problems tend to require more.
+Thus, the early stopping mechanism is triggered when the code population maintains the same fitness value and passes the same test cases for consecutive generations, indicating convergence.
+In Table VIII, we report token consumption for different values of (ranging from to ), denoted as CoCoEvo Stop-1 to CoCoEvo Stop-5, respectively.
+As shown in Table VIII, Sampling consumes the fewest tokens because it only generates code, not test cases.
+Meanwhile, Sampling+Filtering [5], MBR-Exec [13], and CodeT [12] show identical token consumption as they share the same generated programs and test cases, differing only in their final program selection logic.
+Code repair methods (e.g., Self-Repair [7], Reflexion [6], INTERVENOR [8], CodeCoT [12], and AgentCoder [16]) consume significantly more tokens because their prompts must include the original code and test execution results.
+Reflexion consumes the most input tokens because it first generates a reflective analysis before rewriting the code.
+CoCoEvo’s input token usage is higher than CodeT’s because its crossover, mutation, and test generation modules require existing programs and test cases as input prompts.
+However, its output token usage is lower than that of repair-based methods.
+Notably, we observe that with the integration of the early stopping strategy, the token consumption of CoCoEvo is significantly reduced.
+The required input tokens become comparable to those of CodeT, while the output token usage is lower.
+To examine the impact of the early stopping strategy on accuracy, we report the pass@1 performance of CoCoEvo on the LeetCode-Contest dataset under different values of .
+The results are presented in Table IX.
+We find that, when jointly considering token consumption and accuracy, setting the early stopping parameter to leads to a slight decrease in accuracy compared to configurations without early stopping, but achieves a 30% to 50% reduction in overall token usage, demonstrating the considerable potential of CoCoEvo.
+V-H Evaluation on Real-World Projects
+Project | Link | Problems |
+mrjob | https://github.com/Yelp/mrjob | 1 |
+Jinja2 | https://github.com/pallets/jinja | 1 |
+twilio-fatisar | https://github.com/twilio/twilio-python | 1 |
+feedparser | https://github.com/kurtmckee/feedparser | 1 |
+dash | https://github.com/plotly/dash | 2 |
+diffprivlib | https://github.com/IBM/differential-privacy-library | 3 |
+databases | https://github.com/encode/databases | 1 |
+This section evaluates CoCoEvo’s performance on real-world software projects.
+We selected problems from the DevEval dataset [47], where each problem consists of a function and its corresponding unit tests extracted from a real-world project.
+To ensure realistic and manageable problems, we filtered for functions with: (1) at least one external dependency, (2) five or more unit tests, and (3) to lines of code.
+In total, we selected problems from software projects, as summarized in Table X.
+Method | Pass@1 |
+Sampling | 38.00 |
+Sampling+Filtering | 40.00 |
+CodeT | 40.00 |
+CoCoEvo | 44.00 |
+Unlike LeetCode-Contest, these real-world problems have complex dependencies and require full unit tests, not just single assertions.
+To address this, we manually added dependencies to the input prompts, adapted the test generation process to produce full unit tests, and refined the problem descriptions for clarity.
+Using the same setup as Section IV-E, we ran experiments with Qwen2.5-Coder-32B [3] and compared CoCoEvo against three baselines: Sampling, Sampling+Filtering [5], and CodeT [12].
+As shown in Table XI, while all methods’ performance degraded on real-world problems, CoCoEvo still consistently outperformed the baselines.
+A case study in the supplementary material presents a problem solved only by CoCoEvo, where every baseline fails.
+However, our method still exhibits some limitations:
+Dependency Extraction: Dependencies must be manually supplied as the model cannot automatically identify external modules. Future work could automate this with dependency analysis, LLM agents, or RAG.
+Context Limitations: Numerous dependencies can exceed model token limits, requiring memory management or larger context windows.
+Quality of Problem Descriptions: Real-world projects often have unclear descriptions, challenging the LLM’s intent-understanding capabilities.
 VI Conclusion
-This paper introduced CoCoEvo, an LLM-based co-evolution framework that advances automated code generation by integrating simultaneous program and test case evolution.
+This paper introduces CoCoEvo, an LLM-based framework that simultaneously evolves programs and test cases for automated code generation.
 By eliminating the reliance on pre-defined test cases, CoCoEvo addresses a critical limitation in existing approaches, particularly in scenarios lacking comprehensive testing resources.
-The dynamic crossover rate scheduling mechanism and the multi-objective optimization for test case selection were key to the framework’s success, enabling effective exploration and refinement throughout the evolution process.
-Experimental results on the LeetCode-Contest dataset, utilizing four leading LLMs, demonstrated that CoCoEvo achieved superior performance compared to traditional methods.
+The dynamic crossover rate scheduling mechanism and the multi-objective optimization for test case selection are key to the framework’s success, enabling effective exploration and refinement throughout the evolution process.
+Experimental results on the LeetCode-Contest dataset, utilizing four leading LLMs, demonstrate that CoCoEvo achieves superior performance compared to traditional methods.
 Beyond its practical implications in software development, CoCoEvo opens new avenues for integrating evolutionary algorithms with LLMs.
-Future work will focus on enhancing the scalability of CoCoEvo for project-level code generation and incorporating additional quality metrics for test case evaluation.
-Furthermore, exploring its application in other domains, such as automated debugging and software maintenance, could further expand its impact.
+The limitations of CoCoEvo are primarily reflected in its relatively high token consumption.
+However, this can be significantly mitigated by early stopping strategies, which reduce output tokens, though input prompts may remain long.
+Future work could explore how to compress the length of input prompts.
+Moreover, CoCoEvo faces several challenges when applied to real-world software programming tasks.
+Specifically, it lacks precise dependency extraction and memory management capabilities, and it relies heavily on detailed problem descriptions and docstrings.
+Future research may focus on integrating CoCoEvo with techniques such as dependency analysis, LLM-based agents, and RAG, as well as incorporating effective memory management mechanisms, to better adapt the method for real-world software development scenarios.
+In addition, all the datasets used in the experiments of this paper are based on the Python programming language.
+However, CoCoEvo is designed language‐agnostic.
+Future research may investigate the effectiveness of CoCoEvo across multiple programming languages.
+Furthermore, the evolutionary operators in CoCoEvo rely to some extent on the programming capabilities of LLMs.
+Subsequent work could explore how to adapt CoCoEvo for use with smaller or less capable code language models.